The circuit diagram for the PMC servo card is D980352. From this diagram, I see that I can send an excitation from the network analyzer to FP2TEST (9.09 kOhm input impedance) where it is added to the PMC error signal before going to the loop filters.
I hook up the following
I 'Enable' Test 2 on the PSL screen, so FP2TEST gets added to the error signal.
I was able to see the carrier and both sidebands.
I tried to grab this data from the scope via ethernet, but was unsuccessful (timeout errors, I'm using the scripts from scripts/tektronix/tek-dump, and the GPIB box that Kruthi had been using for the GigE cam; I also tried plugging in directly scope->ethernet. Never got anything but timeout errors, so maybe I'm not specifying the port correctly. Anyway the trace is frozen on the scope for later use, or I can easily repeat this now that I know how).
Next, I locked the PMC (Test1 is off, tune DC output adjust until I get some transmission, turn on the loop at Test1, increase the gain to before the loop goes unstable). I'm sending the following channels to SR560 (gain = 2, no filtering, high dynamic reserve, 50 Ohm outputs), and reading spectra from the Agilent 4395A:
The HV mon was always saturating the preamp, so I disconnected it; I added a 50 Hz (6db) high pass to the Trans PD signal, since it has a DC component.
I got to take a look at the traces on the spectrum analyzer front panel, but too tired to do the GPIB for now. There are peaks, things look reasonable.
Today, I tried aligning it further; I'm attaching a picture of it. We are not able to see all the 4 OSEMs yet. In the reference picture I had taken, before taking off the previous analog setup, the OSEMs are not seen. So, I don't really understand what the other 2 spots seen on the current screen are. Are they actually OSEMs?
I need a laptop next to MC2, so that I can have a look at it and make further alignments. So, I tried accessing the GigE attached to the telescope using Paola. The pylon app in it, throws an error, few seconds after running it in continuous shot mode, and disconnects the GigE; everything works fine on Rossa though. I'll put up further details soon.
don't need to lock - make sure the 4 OSEMs are centered on the camera field just as we have for the arm cavity mirrors
I'm attaching a picture of the screen. I just positioned the enclosure by turning it a bit and I suppose we can see the mirror inside the vacuum now (the MC2 is still not locked).
SURFS want some locking of IMC for camera adjustment.
So I left the IMC with intermediate gains so that it keeps locking and unlocking.
VCO (overall) iMC gain of -32, FSS common gain 3, and the FAST gain 20. I believe MC2tickle is ON too.
SURFs had trouble connecting paola to martian via wireless.
Of course, it requires a fixed IP but it had not it yet. So I went to chiara and gave 192.168.113.110 as "paolawl". Note that the wired connection has .111 and it is "paola".
Followed the instruction on http://nodus.ligo.caltech.edu:8080/40m/14121
As directed by Gautam, I have set up one script- interact.py (at /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon/interact.py) to perform the following two tasks:
Steps to view GigE feed for a fixed amount of time:
Steps to record the GigE feed for a fixed amount of time:
I tried to look for elegant solutions that wouldn't require editing the code that Jon wrote and stumbled upon this useful bit of information but ended up deciding that it was just easier to change the camera_client_movie.py (/opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon/camera_client_movie.py). It can still be run as previously described, where video recording is terminated by using Ctrl-C. Steps to record for a fixed period of time are
I'll make aliases for these to make the whole process more user friendly. I'm halting this for now and will discuss what else needs to be done once Gautam gets back.
Regarding the autolocker: I spoke to Aaron today and as he is in tomorrow, I'll ask him about the burt files and the ideal configuration.
I'm also starting with GANs now.
Continuing this investigation of the IMC, today I am getting familiar with the PMC and FSS. I'd like to measure the frequency noise of the PSL referenced to the PMC.
I checked that the PSL shutter is off, so no light reaches the IMC.
I'm not really sure what I'm looking for on the FSS boxes. I found a few documents to guide:
I ran the FSS autolock script from C1PSL_FSS, nothing obvious changes when I do so. The FSS error signal (which I think is PSL-FSS_MIXERM) is flatlined, and the RC-RF_PD has no LO (PSL-FSS_LOCALC is nan).
Rich dropped by at around 3:00 PM today and picked up the VCO in Attachment #1 and left the note in Attachment #2 on Gautam's desk with the promise of bringing it back soon.
Today, Rana asked me to work on improving simulations based on the ideas we discussed last week. As of the previous elog the simulation accomodated only
Today, I added the simulation of point scatterers.
The image on the sensor (camera) is produced in roughly the following steps.
Herewith, in attachments #1, #2, #3 I am attaching videos obtained by varying scattering amplitude and number of scattering points in a vain attempt to reproduce this data. I shall work more on this simulation on Friday.
Neural network stuff:
GANs for simulation:
Networks for beam tracking:
Yesterday, Koji helped me clean all the optics that are being used for the setup. We tried aligning the cameras with the previous configuration we had, but after connecting the analog camera cables there wasn't much room to align the beam splitter. Today, I tried a different configuration and tested the alignment of analog camera, GigE, beam splitter and the mirror using a laser beam [pictures attached]. But the MC2 isn't locked to test if the whole setup is actually aligned with the mirror inside the vacuum.
Also, with this setup, just by using posts of different lengths, we will be able to image the beam spot in all the cavities.
Also, with this setup, just by using posts of different lengths with the middle 90º-post-clamp, we will be able to move all the components. This way, we can easily image the beam spot in all the cavities.
I drew out some idea of how we might use a single OMC to clean both paths of the BHD after mixing, without being susceptible to polarization-dependent effects within the OMC. Basically, can we send the two legs of the BHD into the OMC counterpropagating. I've attached a diagram.
I think one issue would be scattered light, since any backscatter directly couples into the counterpropagating mode, and thus directly to the PD. However, unless the polarization of the scattered light rotates it would not scatter back to the IFO. And, since the LO and signal mix before the OMC, this scattered light would not directly add phase noise.
Maybe more problematic would be that if the rejection at the PBS (or the polarization rotation) isn't perfect, light from the LO directly couples into the dark port. Can we get away with a Faraday isolator before the OMC?
Today, with Milind's help, I installed the analog camera into the MC2 enclosure [picture attached]; but it is not yet focused. We replaced the bulky angular bracket with a simple one, this saved a lot of space inside and it's easier to align other components now. I'll finish setting it up tomorrow.
Telescope design for MC2: Instead of using two 3" long stackable lens tubes (SM2L30), we can use one 3" lens tube with an adjustable lens tube (SM2V10), as shown in the picture. This gives a flexibility to change the focal plane distance by 1" and also reduces the overall length of telescope from 9 inches to 6-7 inches. I decided to use two 150mm biconvex lens instead of a combination of 150mm and 250mm lenses, as the former combination results in lower focal plane distance for a given distance between the lenses.
Specifications of current telescope system (for future reference):
With the above telescope, assuming the MC2 mirror to be at a distance of approx 75cm, the focal plane distance will range from 7.9cm to 8.1cm. Using the adjustable lens tube, we can further make the fine adjustment.
Today, Rana had me key the PSL crate.
Locking the PMC:
Steps to take snapshots using GigE at different exposures [Instructions for Kruthi]:
The python script takes in the above parameters and then takes snapshots by setting the exposure to values starting at minval and going upto maxval incrementing by step at each turn. This uses a simple for loop and is nothing elaborate.
A few unrelated updates:
I managed to fit all the parts into the cylindrical enclosure without having to drill a hole in the enclosure to mount the analog camera (pictures attached); thanks to Koji for helping me find some fancy mechanical components (swivel post clamps, right angle post clamps and brackets). On Thursday, with Chub's help, I took a look at all the current analog camera positions with respect to the cylindrical enclosures. I think this setup gives me enough flexibility to align the components, as necessary, to be able to image the test masses/mirrors in all the cavities. I'll set it up for MC2 tomorrow.
As per Gautam's request, I looked at the IMC situation.
I'll complete the entry later.
I did some more calculations based on our discussions at the meeting yesterday. Posting preliminary results here for comments.
Attachment #1 - Schematic illustration for the scattering scenarios. For all three scenarios, we would like for the scattered field to be lower than unsqueezed vacuum (safety factor to be debated).
Attachment #2 - Requirements on a fraction of the counter-propagating resonant mode of the OMC scattering back into the antisymmetric port, as a function of RIN and phase noise on this field (y-axis) and amount of field (depends on the amount of contrast defect light which can become resonant in the counter propagating mode). I don't encode any frequency dependence here.
Attachment #3 - Requirements on the direct scatter from the arm cavity resonant field (assumed to dominate any contribution from the PRC) onto the OMC DCPDs, for some assumed phase noise (y-axis) and fraction of the field that makes it onto the OMC DCPDs. This is a pretty stringent requirement. But the probability is low (it is the product of three presumably small numbers, (i) probablity of the beam scattering out of the TEM00 mode, (ii) BRDF of the scattering surface, (iii) probability of scattering back towards the DCPDs), so maybe feasible? I didn't model any RIN on this field, which would be an additional noise term to contend with. The range of the y-axis was chosen because I think these are reasonable amplitudes for chamber wall / other scattering surface motion at acoustic frequencies.
Today I ran into the following errors:
Therefore, Koji and I took a look at it and putting our faith in Gautam's hunch from elog 13023, we walked down to rack 1Y1 and keyed it. Following this, all the functionality previously described was restored! Koji then took a look at all the channels handled by this machine and bestowed upon me the permission to key the crate should I lose control of the GigE again.
Thanks! It does indeed do the trick! With that I was able to
Further, a quick look at the camera server code in /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon/camera_server.py revealed that the script expects the details of "Number of Snapshots" in "Camera Settings" in the configuration file i.e in C1-CAM-ETMX.ini at ( /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon/C1-CAM-ETMX.ini) which wasn't present before. Adding this parameter to the config file allows one to take a snapshot using the medm screen. Infact, unlike as described in this elog, I was able to start the server and client as described in elog 14649, and then obtain snapshots using the terminal command caput C1:CAM-ETMX_SNAP 1.
caget/caput probably does the job.
Still not sure about how to modify the exposure time (other than using the pylon app, the only technique I know so far is to adjust the exposure manually on the medm screen and then run the scripts as described in the previous elog).
Figured out how to get/grab frames by looking at the pypylon documenation as that turned out to be easier than modifying Jon's code. Still not sure about how to modify the exposure time (other than using the pylon app, the only technique I know so far is to adjust the exposure manually on the medm screen and then run the scripts as described in the previous elog). I will figure that out tomorrow and make a script suitable for Kruthi's usage (obtain a bunch of images with different exposure times). I will also try and integrate the video saving and streaming code into this and have a neat little script set up asap.
I briefly managed to lock the IMC today - it stayed locked for ~10 minutes. Attachment #1 shows spectra of a few error and control signals for today's lock, and from a stretch yesterday before the problems surfaced*. The 60 Hz lines are much bigger, and MC_F signals broadband excess noise above a few Hz. I suspect a problem somewhere in the electronics.
*I confess the comparison isn't entirely valid because I had to tweak the FSS FAST gain from its nominal value of 22 to 25 in order to get the PC drive RMS down to the ~1.5V level. At the nominal gain setting, with the laser frequency locked to the cavity length, the PC Drive RMS was ~4 V. Still, indicative of something being off in the electronics.
Attachment #1 shows the RIN and phase noise requirements for the 40m BHD for measuring Ponderomotive squeezing.
Chub and I are trying to figure out a way to co-mount GigE into the existing cylindrical enclosure. I'm attaching a picture of the current setup that is being used for imaging MC2. As of now, I have thought of 3 possible setups (schematics attached); but I don't know how feasible they are. Let us know if you have any other ideas.
Update: The setup 3 would require us to use the 52cm long enclosure. It has a long breadboard welded to it, which makes it very convienient, but the whole setup becomes quite heavy and it's not that safe to install such heavy enclosure on top of the vaccuum system. Also, aligning its components would be more complicated than other setups.
I decided to start with the simple one, therefore, I tried implementing setup 1. Fitting in the analog camera horizontally alongside the telescope turned out to be tricky. Though I did manage to fit it in, it didn't leave any room to change the orientation of the beamsplitter. Like Koji suggested, I'll be trying the setup 2.
I was working with the git repo in the SnapPy_pypylon folder (/cvs/cds/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon) and needed to create a branch. To avoid any confusion, I modified the PS1 variable and that alone in the bashrc file to reflect the git branch so that the prompt now displays the git branch if you enter a repository. This is just an update.
The following steps summarize the steps to setting up and interacting with a GigE camera.
Launching the PylonViewerApp:
Using setup python scripts to interact with the GigE (a summary of the steps listed here and here)
Since ~ 2 hours ago, the IMC autolocker has not been able to keep the IMC locked. I don't see any obvious trends in the wall StripTool that may point to what's going on. For the brief periods in which a TEM00 mode is locked, the PC Drive RMS level is ~5x what the nominal level is, and while the autolocker is trying to lock the IMC, the PC drive RMS level is hovering around 4V DC, which is high. The PMC Error and Control signal spectra show huge 60 Hz (and harmonics) peaks, and indeed this is visible in the time domain signals as well (on ndscope or on the oscilloscope on the PSL table), but this is not a new feature in the last two hours. Usually, this kind of problem signals that either/both the c1psl or c1iool0 slow machines need to be power-cycled, but I confirmed that both machines are online and telnet-able. Possibilities: (i) some card in the c1psl / c1ioo crates have failed or (ii) something in the MC/FSS electronics chain has failed or (iii) there is a huge amount of excess high-frequency noise from the NPRO.
I am leaving the PSL shutter closed.
no BMP files
Coupling into the fast axis of the fiber:
The PM couplers I bought require that the light is coupled to the fast axis. The Thorlabs part that Andrew ordered, and which Anjali was using for the MZ experiment, was the opposite configuration, and so the input coupler K6XS mount was rotated to accommodate this polarization. The HWP was also rotated to cut the power into the fiber. I undid these changes. Mode-matching is ~65% (2.42mW/3.70mW) which isn't stellar, but good enough. The PER is ~15dB (ratio of power in fast axis to slow axis is ~40), which I verified using another collimator at the output, and a PBS + two photodiodes. Again isn't stellar but good enough.
EX laser temperature adjustment:
Rana adjusted the temperature of the main laser to 30.61 C. According to the calibration, the EX laser temperature needed to be ~32.8 C. It was ~31.2 C. I made the change by rotating the dial on the front panel of the EX laser controller. Fine adjustment was done using the temperature slider on the ALS screen. With an offset of ~+610 counts, I found a beat at ~80 MHz.
First look at PM beamsplitters:
From my initial test, the beat amplitude was stable to my moving of the fibers . The NF1611 DC monitor reports 2.6 V DC with only the EX light, and 3.15 V DC with only the PSL light. So I should probably cut the PSL power a little to improve the contrast. Assuming the 10 kohm DC transimpedance spec can be believed, this means the expected signal level is 4*sqrt(260uA * 315uA)*700V/A ~0.8 Vpp, and I see ~0.9 Vpp, so roughly things add up (this is actually more consistent with an RF transimpedance of 800V/A, which is maybe not unreasonable). The RF amps for routing this signal to the delay line has been borrowed for the 2um frequency noise experiemnt - I will reacquire it today and check the ALS noise performance.
So overall, I am happy with the performance of the current iteration of the BeatMouth.
Yesterday, we were able to capture some images of objects at a distane of approx 60cm (see the attachment), with the GigE mounted onto the telescope. I think, Johannes had used it earlier to image the ETMX (https://nodus.ligo.caltech.edu:8081/40m/13375). His elog entry doesn't say anything about the focal length of the lenses that he had used. The link to the python code he had used to calculate the lens solution wasn't working. After Gautam fixed it, I took a look at it. He has used 150mm (front lens) and 250mm (back lens) as the focal length of lenses for the calculation. Using the lens formula and an image of a nearby light source, with a very rough measurement, I found the focal lengths to be around 14 cm and 23 cm. So, I'm assuming that the lenses in the telescope are of same focal lengths as in his code, i.e 150mm and 250mm.
To maintain PM fibers all the way through to the photodiode, I had ordered some PM versions of the 50/50 fiber beamsplitters from AFW technologies. They arrived some days ago, and today I installed them in the BeatMouth. Before installation, I checked that the ends of the fibers were clean with the fiber microscope. I also did a little cleanup of the NW corner of the PSL table, where the 1um MZ setup was completely disassembled. We now have 4 non-PM fiber beamsplitters which may be useful for non polarizaiton sensitive applications - they are stored in the glass-door cabinet slightly east of the IY chamber along the Y arm, together with all the other fiber-related hardware.
Anjali had changed the coupling of the beam to the slow axis for her experiment but I ordered beamsplitters which have the slow axis blocked (because that was the original config). I need to revert to this config, and then make a measurement of the ALS noise - if things look good, I'll also patch up the Y arm ALS. We made several changes to the proposed timeline for the summer but I'd like to see this ALS thing through to the end while I still have some momentum before embarking on the BHD project. More to follow later in the eve.
Get a fiber BS that is capable of maintaining the beam polarization all the way through to the beat photodiode. I've asked AFW technologies (the company that made our existing fiber BS parts) if they supply such a device, and Andrew is looking into a similar component from Thorlabs.
Today, we tried to resuscitate the c1iscaux2 channels by swapping the existing, failed VME crate with the newly freed up crate from c1susaux. In summary, the crate gets power, and the EPICS server gets satrted, but I am unable to switch the whitening gain on the whitening boards. I belive that this has to do with the FAIL LEDs that are on for the XVME-220 units. We were careful to preserve the location of the various cards in the VME crates during the swap. Rather than do a detailed debugging with custom RJ45 cables and terminal emulators, I think we should just focus the efforts on getting the Acromag system up and running.
Our work must have bumped a cable to the c1lsc expansion chassis in the same rack - the c1lsc FE had crashed. I rebooted it using the script - everything came back gracefully.
The vacuum itself was fine - CC1 gauge reported a pressure of 1.3e-5 torr. Note to self: the C1:Vac-CC1_HORNET_PRESSURE channel, which is the analog readback of the Hornet gauge and which is hooked up to an Acromag ADC in the c1auxex chassis, is independent of the status of the c1vac machine, and so can serve as a diagnostic.
However, I was unable to interact with c1vac in any way, the monitor hooked up directly to it was showing a frozen display. So I hard-rebooted the system. It took a few minutes to come back online - but even after 10 minutes of waiting, still no display. In the process of the reboot, several valves were closed off - when the EPICS processes restart, there are momentary instances where the readback channels get an "undefined" value, which prompts the main interlock process to transition to a "SAFE" state.
Running df -h, I saw that the /var partition was completely full. Maybe this was somehow interfering with the machine running smoothly? Two files in particular, daemon.log and daemon.log.1 were ~1GB each. The contents of these files seemed to be just the readbacks for the caget and caput commands. So I cleared both these files, and now the /var partition usage is only 26%. I also got the display back up and running on the physical monitor hooked up to the c1vac machine's VGA port. Let's see if this has improved the stability situation. The CPU load is still high (~6-7), with most of this coming from the modbus process. Why is this so high? c1susaux has more Acromag units but claims a much lower load of 0.71. Is the CPU of the c1vac machine somehow inferior?
In the meantime, I ssh-ed into c1vac and restored the "Vacuum normal" valve config. During this little escapade, the main volume pressure rose to ~6e-5 torr. It's coming back down smoothly.
Unrelated to this work: we had turned the RGA off for the vent, I powered it back on and re-initialized it this morning.
I've been monitoring the status of the pumpdown remotely with ndscope lookbacks of C1:Vac-CC1_pressure. Today morning, I saw that the channel was putting out a constant value (signature of EPICS server being frozen). caget did not work either. Then I tried ssh-ing into c1vac to see if there were any issues but I was unable to. The machine isn't responding to ping either. The EPICS value has been frozen since ~1030pm PDT 26 May 2019.
I will try and head to campus later today to check on it. Isn't an email alert or soemthing supposed to be sent out in such an event?
On Friday, I tried calibrating the CCD with the following setup. Here, I present the expected values of scattered power (Ps) at s = 45°, where s is scattering angle (refer figure). The LED box has a hole with an aperture of 5mm and the LED is placed at approximately 7mm from the hole. Thus the aperture angle is 2*tan-1(2.5/7) ≈ 40° approx. Using this, the spot size of the LED light at a distance 'd' was estimated. The width of the LED holder/stand (approx 4") puts a constraint on the lowest possible s. At this lowest possible s, the distance of CCD/Ophir from the screen is given by . This was taken as the imaging distance for other angles also.
In the table below, Pi is taken to be 1.5mW, and Ps and were calculated using the following equations:
Lowest possible s (in degrees)
Expected Ps at s = 45° (in µW)
On measuring the scattered power (Ps) using the ophir power meter, I got values of the same order as that of expected values given the above table. Like Gautam suggested, we could use a photodiode to detect the scattered power as it will offer us better precision or we could calibrate the power meter using the method mentioned in Johannes's post: https://nodus.ligo.caltech.edu:8081/40m/13391.
Things yet to be done:
At ~4pm, the main volume pressure (CC1) was reported to be ~5e-5 torr. So I replaced the HR mirror in the MC REFL path with the usual 10% beamsplitter, and aligned the beam onto MCREFL photodiode. I also replaced the ND filter on the AS port camera, and in front of the IPPOS QPD.
Then I turned up the power by HWP rotation - at the input to the IMC, I now measured 960 mW with the Coherent power meter, so the NPRO power has certainly decayed by ~10% from 2018 July. Normal high-power IMC autolocker script was re-enabled on megatron (and the slow servo enable threshold raised from 1000 cts to 8000cts). IMC was readily locked, after some hand alignment, I got a maximum of 14500 cts transmission. I was then able to lock the Y-arm. The dither alignment servo did not work with the nominal settings, but by hand alignment, I was able to get TRY up to 0.6 (I didn't try too hard to optimize this in any systematic way). X arm was also locked.
AUX drypump valved off and shutdown at ~610pm. I also switched both TP2 and TP3 to their lower rotation "standby" mode. So overall no major mishaps this time around. I am leaving the PSL shutter open over the long weekend. For in-air vs vacuum suspension spectra comparison, I kicked the ETMY optic at Fri May 24 18:26:10 PDT 2019.
Overnight, the pressure of the main volume only rose by 10 mtorr, so there was no need to run the roughing pumps again. So we went straight to the turbos - hooked up the AUX drypump and set it up to back TP2. Initially, we tried having both TP2 and TP3 act as backing pumps for TP1, but the wimpy TP3 current was always passing the interlock threshold. So we decided to pump down with TP3 valved off, only TP2 backing TP1. This went smooth - we had to keep an eye on P2, to make sure it stayed below 1 torr. It took ~ 1 hour to go from 500 mtorr to 100 mtorr, but after that, I could almost immediately open up RV2 completely. A safe setting to run at seems to be to have RV2 open by between 0.5 and 1 turn (out of the full range of 7 turns) until the pressure drops to ~100 mtorr. Then we can crank it open. We are, at the time of writing, at ~8e-5 torr and the pressure is coming down steadily.
I had to manually clear the IG error on the CC1 gauge, and re-enabled the High Voltage, so that we have a readback of the main volume pressure in that range. I made a script to do this (enable the HV, the IG error still has to be cleared by pushing the appropriate buttons on the Hornet), it lives at /opt/target/python/serial/turnHornetON.py. I guess it'll take a few days to hit 8e-6 torr, but I don't see any reason to not leave the turbos running over the weekend.
Remaining tasks are (i) disconnect the roughing pump line and (ii) pump down the annuli, which will be done later today. Both were done at ~2pm, now we are in the vacuum normal config. I'll turn the two small turbos to run on "Standby Mode" before I head home today. I think TP3 may be close to end-of-life - the TP3 current went up to 1A even while evacuating the small volume of the annular line (which was already at 1 torr) with the AUX drypump backing it. The interlock condition is set to trip at 1.2A, and this pump is nominally supposed to be able to back TP1 during the pumpdown of the main volume from 500 mtorr, which it wasn't able to do.
[chub, koji, gautam]
Close up photos of the EY and IY chambers may be found here.
Update on the display manager of c1vac: I was able to get it working again by running sudo systemctl restart display-manager. Now I can interact with the MEDM screens on c1vac. It is a bit annoying that this machine doesn't have the users directory so I don't have access to the many convenient StripTool templates though - maybe I'll make local copies tomorrow for the pumpdown.
On Tuesday, I tried reproducing Pooja's measurements (https://nodus.ligo.caltech.edu:8081/40m/13986). The table below shows the values I got. Pictures of LED circuit, schematic and the setup are attached. The powermeter readings fluctuated quite a bit for input volatges (Vcc) > 8V, therefore, I expect a maximum uncertainity of 50µW to be on a safer side. Though the readings at lower input voltages didn't vary much over time (variation < 2µW), I don't know how relaible the Ophir powermeter is at such low power levels. The optical power output of LED was linear for input voltages 10V to 20V. I'll proceed with the CCD calibration soon.
I have done the following thus far since elog #14626:
I will wrap up the simulation code today and proceed to going through Gabriele's repo. I will also test if the contour detection method works with the simulated data. During our meeting, it was pointed out that when working with real data, care has to be taken to synchronize the data with the video obtained. However, I wish to put off working on that till later in the pipeline as I think it doesn't affect the algorithm being used. I hope that's alright (?).
I did the following:
I think this completes the pre-pumpdown alignment checks we usually do. The detailed plan for tomorrow is here: please have a look and lmk if I missed something.
Yesterday we noticed that the POS and SIDE eigenmodes were degenerate (with 1mHz spectral resolution). Moreover, the YAW peak had shifted down by ~500 mHz compared to earlier this week, although there was still good separation between PIT and YAW in the Oplev error signals. Ideas were (i) check if EQ stops were not backed out sufficiently, and (ii) look for any fibers/other constraints in the system. Today morning, I inspected the optic again. I felt the EQ stop viton tips were a bit close to the optic, so I backed them out further. Apart from this, I adjusted the LR and SIDE OSEM position in their respective holders to make the sensor voltages closer to half-light. Kicked the optic again just now, let's see if there is any change.
If everything goes smoothly, I think we should plan for the heavy doors going back on and commencing the pumpdown tomorrow. After discussion with Koji, we came to the conclusion that it isn't necessary to investigate IPANG (high likelihood of it falling off the steering optics during the pumpdown) / AS beam clipping (no strong evidence that this is a problem) for this vent.
Update 1235: Indeed, the eigenmodes are back to their positions from earlier this week. Indeed, the POS and SIDE modes are actually better separated! So, the OSEM/magnet and EQstop/optic interactions are non-negligible in the analysis of the dynamics of the pendulum.
We executed this plan. Photos are here. Summary:
So if nothing, we got to practise this new wiping technique with OSEMs in situ successfully.
So the primary vent objectives have been achieved, I think.
Tomorrow and later this week:
While we have the chance:
Unrelated to this work: megatron is responding to ping but isn't ssh-able. I also noticed earlier to day that the IMC autolocker blinky wasn't blinking. So it probably requries a hard reboot. I left the lab for tonight so I'll reboot it tomorrow, but no nds data access in the meantime...
For good measure:
Went through all of Pooja's elog posts, her report and am currently cleaning up her code and working on setting up the simulations of spot motion from her work last year. I've also just begun to look at some material sent by Gautam on resonators.
This week, I plan to do the following:
1) Review Gabriele's CNN work for beam spot tracking and get his code running.
2) Since the relation between the angular motion of the optic and beam spot motion can be determined theoretically, I think a neural network is not mandatory for the tracking of beam spot motion. I strongly believe that a more traditional approach such as thresholding, followed by a hough transform ought to do the trick as the contours of the beam spot are circles. I did try a quick and dirty implementation today using opencv and ran into the problem of no detection or detection of spurious circles (the number of which decreased with the increased application of median blur). I will defer a more careful analysis of this until step (1) is done as Gautam has advised.
3) Clean up Pooja's code on beam tracking and obtain the simulated data.
4) Also data like this (https://drive.google.com/file/d/1VbXcPTfC9GH2ttZNWM7Lg0RqD7qiCZuA/view) is incredibly noisy. I will look up some standard techniques for cleaning such data though I'm not sure if the impact of that can be measured until I figure out an algorithm to track the beam spot.
A more interesting question Gautam raised was the validity of using the beam spot motion for detection of angular motion in the presence of other factors such as surface irregularities. Another question is the relevance of using the beam spot motion when the oplevs are already in place. It is not immediately obvious to me how I can ascertain this and I will put more thought into this.
Following the observation that the response in the LL shadow sensor was lower than that of the others, I decided to pull it out a little to move the signal level with nominal DC bias voltage applied was closer to half the open-voltage. I also chose to rotate the SIDE OSEM by ~20 degrees CCW in its holder (viewed from the south side of the EY chamber), to match more closely its position from a photo prior to the haphazhard vent of the summer of 2018. For the SIDE OSEM, the theoretical "best" alignment in order to be insensitive to POS motion is the shadow sensor beam being horizontal - but without some shimming of the OSEM in the holder, I can't get the magnet clear of the teflon inside the OSEM.
While I was inside the chamber, I attempted to minimize the Bounce/Roll mode coupling to the LL and SIDE OSEM channels, by rotating the Coil inside the holder while keeping the shadow sensor voltage at half-light. To monitor the coupling "live", I set up DTT with 0.3 Hz bandwidth and 3 exponentially weighted averages. For the LL coil, I went through pi radians of rotation either side of the equilibrium, but saw no significant change in the coupling - I don't understand why.
In any case, this wasn't the most important objective so I pushed ahead with recovering half-light levels for all the shadow sensors and closed up with the light doors. I kicked the optic again at 1712:14 PDT, let's see what the matrix looks like now.
before starting this work, i had to key the unresponsive c1auxey VME crate.